### v1

# The TRGB--SBF Project. III. Refining the HST Surface Brightness Fluctuation Distance Scale Calibration with JWST

Abstract
The TRGB--SBF Project team is developing an independent distance ladder using a geometrical calibration of the tip of the red giant branch (TRGB) method in elliptical galaxies that can in turn be used to set the surface brightness fluctuation (SBF) distance scale independent of Cepheid variables and Type Ia supernovae. In this paper, we use JWST TRGB distances calibrated using the megamaser galaxy NGC 4258 to determine a new Cepheid-independent SBF zero point for HST. This new calibration, along with improved optical color measurements from PanSTARRS and DECam, gives an updated value of the Hubble-Lemaître constant $H_{0}=73.8\pm 0.7\pm 2.3$ km s-1 Mpc-1 that is virtually identical to the previously published result based on SBF distances calibrated using Cepheids (Blakeslee et al. 2021). Future improvements in the geometrical calibration of TRGB and additions to the HST SBF survey will further reduce the random and systematic uncertainties on $H_{0}$ measured using the extensive HST WFC3/IR SBF dataset. JWST observations of the Coma cluster will soon establish a foundation for the SBF calibration that will extend to much larger distances than is possible with HST.
Distance indicators (394) — Galaxy distances (590) — Hubble-Lemaître constant (758)
journal: ApJfacilities: HST (WFC3, ACS), JWST (NIRCam), PS1, Blanco (DECam)
1 Introduction
The current disagreement between the measured expansion rate of the universe and that computed from the standard $\Lambda$CDM cosmological model tied to the cosmic microwave background fluctuations is now greater than 5-$\sigma$ (Planck Collaboration et al., 2020; Riess et al., 2022). Resolving this discrepancy is the most pressing issue in cosmology today. It is not yet clear if the explanation is an accumulation of various systematic errors in local distance measurements or if new physics beyond the $\Lambda$CDM model is required (e.g., Di Valentino et al., 2021).
The TRGB--SBF Project seeks to reduce or eliminate systematic uncertainties in the first rungs of the distance ladder by developing a new distance scale based on a geometrical calibration of old, metal-rich populations in massive elliptical galaxies (i.e., Pop II RGB stars) that is completely independent of the Cepheid and Type Ia Supernova distance measurements. To do this, we are connecting Tip of the Red Giant Branch (TRGB) and Surface Brightness Fluctuation (SBF) distances measured in 14 nearby elliptical galaxies using the James Webb Space Telescope (Tully et al., 2023). The TRGB--SBF Project will eventually link geometrically-calibrated TRGB distances at 15-20 Mpc with SBF distances reaching out to perhaps 300 Mpc or more using the same telescope (JWST), camera (NIRCam), and filters to minimize systematic uncertainties for each rung of the ladder.
SBF distances have already been measured in more than 370 galaxies out to about 100 Mpc using the Hubble Space Telescope (HST), including more than 220 in F110W using the WFC3/IR camera. In this paper, we use the TRGB distances to eight individual galaxies calibrated using the geometrical distance to the megamaser galaxy NGC 4258 (Reid et al., 2019; Anand et al., 2024a) to calibrate the HST SBF distances via TRGB. In the future, the Gaia satellite (Gaia Collaboration et al., 2017, 2018) will add additional geometrical anchors to the TRGB distance scale to further reduce the systematic calibration uncertainty. This study presents the first step towards an SBF calibration that is independent of the Cepheid distance scale and connects JWST TRGB to the existing HST SBF distance measurements out to 100 Mpc.
2 Infrared Surface Brightness Fluctuations with HST
[…]
The current generation of optical and IR HST SBF measurements are based on the calibrations published by Mei et al. (2007); Blakeslee et al. (2009, 2010) and Jensen et al. (2015, 2021). The foundation of all the HST SBF measurements to-date is the HST Key Project Cepheid distances to seven galaxies with ground-based SBF distances measured by Tonry et al. (2001) and Blakeslee et al. (2002), combined with a large number of overlapping ground-based and HST SBF measurements in the Virgo and Fornax galaxy clusters (see Blakeslee et al., 2010, for direct comparisons). These large surveys provided the statistical foundation for SBF and showed that the intrinsic scatter in the SBF technique is  ${\sim\,}0.06$ mag (optical) to  ${\sim\,}0.1$ mag (near-IR) after correction for stellar population variations using optical colors (Blakeslee et al., 2009; Jensen et al., 2015).
Fluctuations are much brighter at near-IR wavelengths than in the optical and are measurable to ${\gtrsim}80$ Mpc in a single HST orbit. SBF is much more efficient than Cepheids or supernovae for distances out to 100 Mpc and achieves comparable precision. The majority of the SBF distances beyond 40 Mpc useful for measuring $H_{0}$ were made using NICMOS (Jensen et al., 2001) and WFC3/IR (Cantiello et al., 2018; Blakeslee et al., 2021). The most recent WFC3/IR F110W SBF calibration was determined using 11 galaxies in the Fornax and Virgo galaxy clusters, and was corrected for stellar population variations using ACS $(g_{475}{-}z_{850})$ colors (Jensen et al., 2015). Updates to the Jensen et al. (2015) calibration to include better Galactic foreground extinction corrections and an improved distance to the LMC were made as described by Cantiello et al. (2018) and Jensen et al. (2021).
The majority of the HST IR SBF observations published between 2015 and 2021 were part of HST programs GO-14219 (PI J. Blakeslee), GO-14654 (PI P. Milne), and GO-15625 (PI J. Blakeslee). Jensen et al. (2021) published SBF distances to 63 galaxies from these programs, including 23 supernova host galaxies. Three more HST programs currently nearing completion, GO-16262 (PI R. Tully), GO-17436 (PI J. Jensen), and GO-17446 (PI P. Milne) have already observed ${\gtrsim}164$ more galaxies out to ${\gtrsim}85$ Mpc, a sample that more than doubles the number of type Ia supernova host galaxies with high-quality SBF distances (now 51). To make the most of these new datasets, in this paper we update the HST SBF calibration using the new JWST TRGB distances.
The SBF distances published by Jensen et al. (2021) were used by Blakeslee et al. (2021) to measure a value of the Hubble-Lemaître constant of $H_{0}=73.3\pm 0.7$ (statistical) ±$\pm 2.4$ (systematic)  km s-1 Mpc-1, using both Cepheid and TRGB calibrations to help reduce systematic uncertainties. The subsample of supernova host galaxies provided an alternate connection between Cepheids and Ia SN for massive early-type galaxies via SBF: Garnavich et al. (2023) measured a value of $H_{0}=74.6\pm 0.9\pm 2.7$  km s-1 Mpc-1 using SBF distances and SN light curve measurements from Scolnic et al. (2022). When the SALT2 fit parameters were re-derived by Garnavich et al. (2023) for the fast-declining SNe typically found in the early-type galaxies, the value of $H_{0}$ became $73.3\pm 1.0\pm 2.7$ km s-1 Mpc-1.
[…]
3 Re-calibrating the HST WFC3/IR SBF Distance Scale using TRGB
JWST observations of the TRGB in 14 nearby elliptical galaxies (GO-3055; Tully et al. 2023) provide a new and direct calibration of IR SBF based on the geometrical megamaser distance to NGC 4258 (Anand et al., 2024a; Reid et al., 2019). We will eventually use JWST to determine $H_{0}$ directly by measuring SBF distances out to  ${\sim}\,300$ Mpc, but building a database of distant SBF measurements will take time, and may never reach the quantities already observed with HST over the last 30 years (more than $370$ galaxies, including SBF distances from NICMOS, ACS, and WFC3/IR). To leverage the full potential of the HST database, we need to connect the new JWST TRGB distances to the existing HST SBF dataset. Three of the JWST TRGB target galaxies in the Fornax cluster and five in Virgo have also been observed with WFC3/IR, as shown in Table 1. The new JWST TRGB measurements of these galaxies (Anand et al., 2024b, 2025) provide us with an opportunity to reduce the systematic uncertainties in the SBF distances originating in the Cepheid calibration. Using a direct geometrical calibration of TRGB from NGC 4258 (Reid et al., 2019; Anand et al., 2024a) and SBF measurements in the same galaxies, using the same telescope, camera, and filters, we can eliminate some sources of uncertainty (e.g., Cepheid metallicity corrections in different galaxy types and positional offsets between spirals and ellipticals in the calibration clusters).
3.1 Updated Calibration
The most direct re-calibration of the HST IR SBF distance scale is to replace the distances from Jensen et al. (2015) with the new JWST-based TRGB distances to the eight galaxies in common between the HST and JWST SBF calibration data sets. The SBF calibration of absolute fluctuation magnitude $\overline{M}_{110}$ depends on the distance scale zero point and a slope with galaxy color, which corrects $\overline{M}_{110}$ for variations in stellar population age and metallicity (e.g., Jensen et al., 2003). Jensen et al. (2015) found that the calibration scatter was lowest when using the average distance to the cluster instead of individual optical SBF distances. We can do the same here, and by using the mean cluster distances from the JWST TRGB measurements we include four more galaxies in Fornax and two more in Virgo, for a total of 14 galaxies in our new calibration (see Table 1).
Table 1: SBF Calibration Data
Galaxy $(m{-}M)_{\rm TRGB}$footnotemark: TRGB $\sigma_{\rm tot}$footnotemark: $(m{-}M)_{\rm clust}$footnotemark: $\overline{m}_{110}$,0footnotemark: DECam $(g{-}z)_{0}$footnotemark: PS $(g{-}z)_{0}$footnotemark: HST
(mag) (mag) (mag) (AB mag) (AB mag) (AB mag) Program
(1) (2) (3) (4) (4, 5) (4) (6)
IC 2006  $\cdots$  $\cdots$ $31.424\pm 0.057$ $28.65\pm 0.021$ $1.274\pm 0.014$  $\cdots$ 11712
NGC 1344  $\cdots$  $\cdots$ $31.424\pm 0.057$ $28.46\pm 0.019$ $1.201\pm 0.012$  $\cdots$ 11712
NGC 1374  $\cdots$  $\cdots$ $31.424\pm 0.057$ $28.58\pm 0.025$ $1.226\pm 0.012$  $\cdots$ 11712
NGC 1375  $\cdots$  $\cdots$ $31.424\pm 0.057$ $28.26\pm 0.021$ $1.075\pm 0.012$  $\cdots$ 11712
NGC 1380 $31.397\pm 0.034$ 0.072 $31.424\pm 0.057$ $28.60\pm 0.022$ $1.283\pm 0.012$  $\cdots$ 11712
NGC 1399 $31.511\pm 0.036$ 0.073 $31.424\pm 0.057$ $28.84\pm 0.023$ $1.363\pm 0.012$  $\cdots$ 11712
NGC 1404 $31.364\pm 0.034$ 0.072 $31.424\pm 0.057$ $28.73\pm 0.024$ $1.325\pm 0.012$  $\cdots$ 11712
NGC 4458  $\cdots$  $\cdots$ $31.055\pm 0.086$ $28.05\pm 0.023$ $1.149\pm 0.012$ $1.170\pm 0.023$ 11712
NGC 4472 (M49) $31.091\pm 0.032$ 0.071 $31.055\pm 0.086$ $28.43\pm 0.021$ $1.383\pm 0.012$ $1.380\pm 0.011$ 11712
NGC 4489  $\cdots$  $\cdots$ $31.055\pm 0.086$ $27.87\pm 0.019$ $1.126\pm 0.013$ $1.109\pm 0.024$ 11712
NGC 4552 (M89) $30.933\pm 0.041$ 0.075 $31.055\pm 0.086$ $28.31\pm 0.022$ $1.349\pm 0.013$ $1.366\pm 0.012$ 15082
NGC 4636 $31.120\pm 0.035$ 0.072 $31.055\pm 0.086$ $28.36\pm 0.021$ $1.332\pm 0.013$ $1.330\pm 0.013$ 17446
NGC 4649 (M60) $31.061\pm 0.034$ 0.072 $31.055\pm 0.086$ $28.52\pm 0.021$ $1.426\pm 0.013$ $1.423\pm 0.011$ 11712
NGC 4697 $30.330\pm 0.036$ 0.073 $30.330\pm 0.073$ $27.52\pm 0.021$ $1.271\pm 0.013$ $1.271\pm 0.011$ 15226
footnotetext: (1) JWST TRGB distances from Anand et al. (2024b) and Anand et al. (2025). JWST data are available through MAST, https://doi.org/10.17909/z9ch-wk24 (catalog DOI: 10.17909/z9ch-wk24).
footnotetext: (2) Total uncertainty for the TRGB distance including systematic uncertainty.
footnotetext: (3) Average cluster TRGB distance moduli for Virgo (including NGC 4636) and Fornax, except for NGC 4697, which is in the foreground (Anand et al., 2024b, 2025). The uncertainties are the cluster depths from the ACS Virgo and Fornax Cluster surveys (Blakeslee et al., 2009), 0.085 and 0.053 mag respectively, added in quadrature with the uncertainty in the cluster averages of the TRGB distances (0.02 mag for the three TRGB galaxies in Fornax and 0.013 mag for the seven Virgo).
footnotetext: (4) SBF magnitudes and optical colors have been corrected for Galactic extinction using Schlafly & Finkbeiner (2011) and the NED extinction calculator https://doi.org/10.26132/NED5 (catalog DOI: 10.26132/NED5). All fluctuation magnitudes and colors were measured in two annular region around each galaxy from 8.2 to 16.4 arcsec and 16.4 to 32.8 arcsec in radius. The values shown are the weighted averages of the two measurements.
footnotetext: (5) DECam $(g{-}z)_{0}$ colors have been transformed to the PanSTARRS photometric system using Equation 1 (Dey et al., 2019) to simplify comparison.
footnotetext: (6) WFC3/IR data for these galaxies are available through MAST, https://doi.org/10.17909/303n-6g35 (catalog DOI: 10.17909/303n-6g35).
The direct overlap between JWST and HST calibrations includes three galaxies in the Fornax cluster and five in Virgo (including NGC 4636 and NGC 4697 in peripheral in-falling regions). Using the combined average distance to Virgo from TRGB (Anand et al., 2024b, 2025) and the ACSVCS survey average distance modulus of $31.092\pm 0.013$ mag (Mei et al., 2007; Blakeslee et al., 2009), which was used to calibrate the previous IR SBF measurements, the weighted mean difference is $\Delta\mu_{TRGB-SBF}=-0.067\pm 0.031$ mag (Anand et al., 2025). For Fornax, the difference between the mean ACS SBF distance modulus of 31.51 mag (Blakeslee et al., 2009) used for the Jensen et al. (2015) IR SBF calibration and the mean distance from three JWST TRGB measurements gives a difference of $\Delta\mu=-0.086\pm 0.077$ mag, where the error results from the scatter among the three galaxies.
[…]
The SBF measurements were updated since the original calibration published by Jensen et al. (2015), and for this study we have included three additional calibration galaxies (NGC 4552, NGC 4636, and NGC 4649). The new $\overline{m}_{110}$ measurements for the calibrators are consistent with the procedures we used to measure $\overline{m}_{110}$ in the sample from which we derived $H_{0}$ (Jensen et al., 2021; Blakeslee et al., 2021). The average difference between the SBF magnitudes $\overline{m}_{110}$ from the Jensen et al. (2015) calibration paper and the present values for the five calibrator galaxies in common is 0.072 mag with a standard deviation of 0.03 mag. There have been a number of procedural improvements to the SBF analysis that were previously applied to the sample used to determine $H_{0}$ (Blakeslee et al., 2021), and those have been applied to these new measurements. We previously identified an offset of $+0.05\pm 0.02$ mag to the 2015 calibration values due to the PSF normalization (Cantiello et al., 2018). Given the small sample size in this study, the 0.072 mag offset is considered to be statistically consistent with the previously measured 0.05 mag offset. Further assessment of the reliability of the $\overline{m}_{110}$ measurements will require additional overlap between calibration samples.
Optical galaxy colors are critical for the SBF method as they are necessary to correct the SBF amplitudes for intrinsic variations in the ages and metallicities of the stellar populations in elliptical galaxies (Jensen et al., 2003; Blakeslee et al., 2009; Jensen et al., 2015). The original IR SBF calibration was based on the ACS $(g_{475}{-}z_{850})$ colors, which are not generally available for the vast majority of the WFC3/IR observations in the HST archive. Jensen et al. (2021) transformed PanSTARRS $(g{-}z)$ colors (Chambers et al., 2019) to the ACS photometric frame for computing distances. For this study, we remeasured all the PanSTARRS $(g{-}z)$ colors and determined the SBF calibration directly in that system. We found that for galaxies with significant color gradients, the PanSTARRS $(g{-}z)$ colors used by Jensen et al. (2021) for computing SBF distances—measured in the high-$S/N$ ratio central regions—did not always match the colors in the regions where the SBF was measured. For typical massive galaxies with slightly redder cores and bluer halos, correcting the colors to match the SBF regions leads to larger distances, by as much as 5% in some cases, and hence this revision tends to reduce the value of $H_{0}$. For the 61 galaxies in the Jensen et al. (2021) dataset with PanSTARRS photometry, which are used again in this paper to recompute $H_{0}$, the color shift was measured to be $-0.023\pm 0.019$ mag (bluer). The average photometric uncertainty in $(g{-}z)$ is 0.02 mag for PanSTARRS, which includes the uncertainty in the extinction correction (Schlafly & Finkbeiner, 2011) added in quadrature. Given that the historical slope of the relationship between SBF $\overline{M}_{110}$ and $(g{-}z)$ is  ${\sim\,}2$ (Jensen et al., 2015), the expected shift in distance modulus is approximately 0.05 mag or 2.3% in distance, counteracting in part the increase in $H_{0}$ expected from the new TRGB zero point.
The original SBF calibration of Jensen et al. (2015) included equal numbers of Fornax and Virgo cluster galaxies, which were calibrated using ACS optical photometry. The JWST TRGB sample includes three galaxies in the Fornax cluster. The PanSTARRS-1 survey does not reach far enough south to include Fornax, so the current TRGB calibration is limited to five galaxies in Virgo. DECam survey provides a higher-$S/N$ alternative to the PanSTARRS calibration in both Fornax and Virgo. A number of imaging surveys (e.g., DESI and DECaLS) using the DECam wide field imager on the Blanco Telescope at CTIO have been photometrically calibrated and are now publicly available as part of the Legacy Surveys (Dey et al., 2019) and can be explored through NOIRLab’s Astro Data Lab. [^1] https://datalab.noirlab.edu/sia.php We retrieved data in the Legacy Survey DR10 release [^2] https://www.legacysurvey.org/dr10/description/ through the server at the National Energy Research Scientific Computing Center (NESRC). We combined the various archival images (“bricks”) for 23 target galaxies in the Jensen et al. (2021) sample and measured their $(g{-}z)$ colors. Uncertainties were determined by similarly combining the variance frames. The DECam Legacy Survey photometric uncertainties are 0.0073 mag for $g$ and $z$, which were combined in quadrature with each other and with the uncertainty in the foreground extinction correction from Schlafly & Finkbeiner (2011). The mean uncertainty in $(g{-}z)$ is 0.011 mag.
Dey et al. (2019) provide a transformation between PanSTARRS and DECam $(g{-}z)$ values as follows:
$\begin{split}(g{-}z)_{\rm PS}=(g{-}z)_{\rm DECam}+0.02521-0.11294(g{-}i)+\\ +0.01796(g{-}i)^{2}-0.00285(g{-}i)^{3}\end{split}$ (1)
where $(g{-}i)$ is the same in both photometric systems. For the seven galaxies with $(g{-}z)$ colors in both systems, the predicted vs. computed values of $(g{-}z)$ differ by $0.002\pm 0.013$ mag (Fig. 1).
Figure 1: Comparison of the PanSTARRS $(g{-}z)$ colors with those transformed from the DECam colors to PanSTARRS using Dey et al. (2019) for the seven calibrator galaxies with color data in both surveys. The black line denotes the one-to-one relation and is not a fit to the data.
Since DECam $(g{-}z)$ colors were measured for both the Fornax and Virgo cluster samples, we used the DECam colors for the calibration, but transposed to the PanSTARRS photometric system because PanSTARRS was used for the published galaxy sample used to measure $H_{0}$ (Jensen et al., 2021; Blakeslee et al., 2021). The values listed in Table 1 are in the PanSTARRS photometric system and have an additional uncertainty of 0.013 mag added in quadrature to account for the observed scatter in the transformation.
The eight galaxies for which JWST TRGB distances were individually measured (black points in the right half of Figure 2) provide a zero point good to 0.018 mag but a relatively weak constraint on the slope due to the limited range in color covered by the JWST sample (the slope is $1.7\pm 0.3$ in the PanSTARRS $(g{-}z)$ photometric system). To determine a more robust measurement of the slope, we adopted the mean cluster TRGB distance for the Virgo and Fornax clusters, and used those distances to compute $\overline{M}_{110}$ for a larger sample of HST Virgo and Fornax calibrators from Jensen et al. (2015) that include a wider range of galaxy color. When we include six additional galaxies and calculate $\overline{M}_{110}$ using mean TRGB cluster distances (see Table 1) we get a tighter value for the slope of $1.86\pm 0.16$ with a reduced $\chi_{\nu}^{2}=0.74$ per degree of freedom. The fact that the reduced $\chi_{\nu}^{2}$ is less than one suggests that the uncertainties in the absolute magnitude may be slightly overestimated, but there is still a 30% probability of measuring this value of $\chi_{\nu}^{2}$ for the number of degrees of freedom in our sample, indicating general consistency. Regardless of which TRGB distances (and slope) are used, the zero point at a color of $(g{-}z)=1.30$ is consistently $-2.76$ with an uncertainty below 0.02 mag.
The calibration calculated using the mean TRGB cluster distances (Fig. 2) is:
$\overline{M}_{110}=(1.86\pm 0.16)[(g{-}z)_{\rm PS}{-}1.30]+(-2.760{\pm}0.016)$ (2)
where the color is measured in the PanSTARRS $(g{-}z)$ photometric system. The calibration in the original DECam $(g{-}z)$ system is:
$\overline{M}_{110}=(1.75\pm 0.14)[(g{-}z)_{\rm DC}{-}1.40]+(-2.728{\pm}0.017).$ (3)
Figure 2: Absolute fluctuation magnitude $\overline{M}_{110}$ as a function of DECam galaxy color converted to the PanSTARRS photometric system for the eight galaxies with TRGB distances (Anand et al., 2024b, 2025) plus six additional galaxies in the Fornax and Virgo clusters previously used to calibrate F110W SBF (blue points). The fit shown was performed using the “emcee” package in Python, iteratively accounting for uncertainties in both axes. The blue line is the best fit to the blue points; the blue envelopes mark the 1-$\sigma$ and 2-$\sigma$ uncertainties in slope and intercept. The absolute magnitudes calculated using the individual TRGB distance moduli are also plotted for five of the galaxies with PanSTARRS colors (green points) and eight galaxies with DECam colors (black points) to show that slopes derived from the individual and cluster distances are consistent.
3.2 Updated SBF Distances
To assess the impact of the new calibration on the IR SBF distance scale, we applied the new TRGB calibration and PanSTARRS colors to the published fluctuation magnitudes for 61 galaxies from Jensen et al. (2021) and Blakeslee et al. (2021), and recomputed their distances. We also measured DECam $(g{-}z)$ colors for 23 galaxies in the distant sample (the rest were too far north to be included in the DECam Legacy Surveys). The new TRGB zero point and updated colors have opposing effects on the measured distances, and the new calibration implies almost no change in the SBF distances from Jensen et al. (2021) or the resulting values of $H_{0}$ from Blakeslee et al. (2021) or Garnavich et al. (2023).
Overall, the new calibration results in an average offset in the distance modulus of $0.011\pm 0.005$ mag with RMS scatter of 0.04 mag. This implies a change in the distance scale of a factor of $0.995\pm 0.002$ relative to the 2021 distances, meaning that our new distances are on average $0.5\%\pm 0.2\%$ closer, and, consequently, the value of $H_{0}$ would be 0.5% larger. We obtain very similar results if we choose to use only the eight calibrators for which direct TRGB distances are available (distance scale factor of $1.000\pm 0.003$) or the five Virgo calibrators with PanSTARRS colors (scale factor $1.007\pm 0.002$), even though the slope of the $\overline{M}_{110}$–color relation is not as well constrained. The result is robust to the selection of galaxy colors (DECam or PanSTARRS) and to the choice of individual or cluster TRGB distances. For example, if we use only DECam colors with no translation to PanSTARRS, and only use the subset of 23 galaxies in the $H_{0}$ sample with DECam colors, we get a mean distance modulus offset of $0.002\pm 0.013$ mag, or a distance scale factor of $0.999\pm 0.006$.
We are confident that the result of applying the new calibration based on the JWST TRGB measurements, including improved optical color measurements, does not change the conclusions of Blakeslee et al. (2021) regarding the value of $H_{0}$ in any significant way. The resulting value of $H_{0}$ using the updated zero point and the published calibration slope gives $73.81$ km s-1 Mpc-1 instead of $73.44\pm 0.7$ km s-1 Mpc-1, the value found in our 2021 work when using only the Cepheid calibration. Since we are only changing the calibration and not the input $\overline{m}_{110}$ fluctuation magnitudes or velocities, the statistical error in the value of $H_{0}$ is unchanged. A similar 0.5% increase would result for the other values of $H_{0}$ calculated from IR SBF distances (Garnavich et al., 2023; Blakeslee et al., 2021).
The reassessment of the calibration was completed independently without reference to the resulting distances to avoid any confirmation bias in the value of $H_{0}$.
3.3 Systematic Uncertainties
The SBF distance scale is primarily limited by systematic uncertainties. The Cepheid calibration of SBF injected a 0.09 mag systematic uncertainty (Blakeslee et al., 2021) that we remove here and replace with that of the NGC 4258 megamaser and JWST TRGB calibration. The systematic uncertainty associated with the maser distance to NGC 4258 is 1.48% (0.032 mag, Reid et al. 2019). The TRGB measurement uncertainty of 0.033 mag for NGC 4258 from Anand et al. (2024a) acts as an additional systematic error in the TRGB distances of other galaxies that are tied to it. This source of systematic error cannot be reduced further without observations of additional fields around NGC 4258 or other geometrical anchors besides NGC 4258. The link from TRGB to SBF contains the zero point uncertainty of 0.016 mag from Equation 2 above. Combined in quadrature, these add up to 0.049 mag. The intrinsic scatter in the properties of the TRGB feature (and how it is measured) has not been fully quantified. Anand et al. (2024a) estimated a scatter of 0.02 mag in their study, and estimates from Riess et al. (2024) for a sample of spiral galaxies observed in multiple filters with JWST suggest larger values from 0.01 to 0.08 mag. The elliptical galaxies we observed span a smaller range in age and metallicity and should have modest intrinsic scatter; we conservatively include another 0.04 mag in our error budget.
Taken together, these terms give a total systematic uncertainty on the TRGB calibration of HST SBF of 0.063 mag, or 2.9% in distance. Combining this with the estimated 1% uncertainty in the overall velocity flow of the volume covered by our sample (see Blakeslee et al., 2021), the total systematic uncertainty in $H_{0}$ becomes 3.1%, or 2.3 km s-1 Mpc-1 in $H_{0}$. This is a significant improvement over the previous Cepheid-based calibration of 0.09 mag and will improve further with additional geometrical anchors tied to TRGB with Gaia and with more links between JWST TRGB and HST SBF.
4 Conclusions
[…]
This updated calibration is a first step in establishing a new distance ladder with JWST that will be independent of the traditional rungs that use the LMC distance, Cepheids and type Ia supernovae. We are optimistic that additional galaxies will be observed by both JWST and HST, increasing the overlap between the two telescopes both for TRGB and SBF measurements, and further reducing systematic uncertainties.
JWST imaging observations of 39 targets in the Coma cluster (likely including  ${\sim\,}50$ or more elliptical galaxies) have been scheduled for Cycle 3 (GO-5989; Jensen et al. 2024) and will provide a robust determination of the calibration slope in a cluster that is foundational to the distance scale. Additional geometrical anchors securing the absolute magnitude of the TRGB will further enhance the precision of the SBF method. The combination of HST and JWST SBF measurements will create a new distance ladder based solely on old, metal-rich populations in early-type galaxies and free from many of the uncertainties in the Cepheid+SN distance ladder.
Acknowledgments. This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. The results presented here used data associated with HST programs GO-11711, 11712, 12450, 14219, 14654, 15265, 14771, 14804, and 15329. J. Jensen and M. Cowles acknowledge partial support from STScI under grants HST-GO-16262 and JWST-GO-03055. This work also used observations made with the NASA/ESA/CSA James Webb Space Telescope program GO-3055. J. Blakeslee is supported by NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation. M. Cowles acknowledges support from The Barry Goldwater Scholarship and Excellence in Education Foundation and the National Science Foundation. M. Cantiello acknowledges support from the project “INAF Exploration of Diffuse Galaxies with Euclid” (INAF-Large Grant 12-2022, P.I. Leslie K. Hunt), and by Italian Ministry for Education University and Research (MIUR) grant PRIN 2022 2022383WFT “SUNRISE,” CUP C53D23000850006 (P.I. Enrica Iodice). G. S. Anand acknowledges financial support from JWST GO–3055.
This project used data obtained with the Dark Energy Camera (DECam), which was constructed by the Dark Energy Survey (DES) collaboration. Funding for the DES Projects has been provided by the DOE and NSF (USA), MISE (Spain), STFC (UK), HEFCE (UK), NCSA (UIUC), KICP (U. Chicago), CCAPP (Ohio State), MIFPA (Texas A&M), CNPQ, FAPERJ, FINEP (Brazil), MINECO (Spain), DFG (Germany) and the Collaborating Institutions in the Dark Energy Survey, which are Argonne Lab, UC Santa Cruz, University of Cambridge, CIEMAT-Madrid, University of Chicago, University College London, DES-Brazil Consortium, University of Edinburgh, ETH Zürich, Fermilab, University of Illinois, ICE (IEEC-CSIC), IFAE Barcelona, Lawrence Berkeley Lab, LMU München and the associated Excellence Cluster Universe, University of Michigan, NSF NOIRLab, University of Nottingham, Ohio State University, OzDES Membership Consortium, University of Pennsylvania, University of Portsmouth, SLAC National Lab, Stanford University, University of Sussex, and Texas A&M University.
This research uses services or data provided by the Astro Data Lab, which is part of the Community Science and Data Center (CSDC) Program of NSF NOIRLab. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA), Inc. under a cooperative agreement with the U.S. National Science Foundation.
The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation.
This research used the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.
This research made use of Montage, which is funded by the National Science Foundation under Grant Number ACI-1440620, and was previously funded by the National Aeronautics and Space Administration’s Earth Science Technology Office, Computation Technologies Project, under Cooperative Agreement Number NCC5-626 between NASA and the California Institute of Technology.